OpenStudy (anonymous):

Simplify 3 over square root of 11

OpenStudy (anonymous):

\[\frac{3}{\sqrt {11}} =\frac{3}{\sqrt {11}} \times \frac{\sqrt {11}}{\sqrt {11}} \] \[= \frac{3\sqrt {11}}{11} \]

OpenStudy (anonymous):

(3srqrt11)/11

OpenStudy (anonymous):

11/11